# Networked Communication for Mean-Field Games with Function Approximation and Empirical Mean-Field Estimation

### v2

@@ 2.3 Networks
Definition 6 (Time-varying network).
@@
We also define an alternative version of the observation graph that is useful in a specific subclass of environments, which can most intuitively be thought of as those where agents’ states are positions in physical space. When this is the case, we usually think of agents’ ability to observe each other as depending more abstractly on whether states are visible to each other. We define this visibility graph as follows:
Definition 7 (Time-varying state-visibility graph).
[…]
We view an agent in $s$ as able to obtain a count of the number of agents in $s^{\prime}$ if $s^{\prime}$ is visible to $s$. The benefit of this graph $\mathcal{G}^{vis}_{t}$ over $\mathcal{G}^{obs}_{t}$ is that there is mutual exclusivity: either an agent in state $s$ is able to obtain a total count of all of the agents in state $s^{\prime}$ (if $s^{\prime}$ is visible to $s$), or it cannot obtain information about any agent in state $s^{\prime}$ (if those states are not visible to each other). Additionally, this graph permits an agent in state $s$ to observe that there are no agents in state $s^{\prime}$ as long as $s^{\prime}$ is visible to $s$. These benefits are not available if the observability graph is defined strictly between agents as in $\mathcal{G}^{obs}_{t}$, such that using $\mathcal{G}^{vis}_{t}$ facilitates more efficient estimation of the global mean-field distribution from local information in settings where $\mathcal{G}^{vis}_{t}$ is applicable (see Sec. 4).
Algorithm 1 Networked learning with non-linear function approximation
0:  loop parameters $K,M,L,E,C_{p}$, learning parameters $\gamma,\tau_{q},|B|,cl,\nu$, $\{\tau^{comm}_{k}\}_{k\in\{0,\dots,K-1\}}$
0:  initial states {$s^{i}_{0}$}${}_{i=1}^{N}$; $t\leftarrow 0$
1:  $\forall i$ : Randomly initialise parameters $\theta^{i}_{0}$ of Q-networks $\check{Q}_{\theta^{i}_{0}}(o,\cdot)$, and set $\pi^{i}_{0}(a|o)=$ softmax$\left(\frac{1}{\tau_{q}}\check{Q}_{\theta^{i}_{0}}(o,\cdot)\right)(a)$
@@
10:        Update $\theta$ to minimise ^$\hat{\mathcal{L}}(\theta,\theta^{\prime})$ as in Def. 8
@@
14:     $\forall i$ : $\pi^{i}_{k+1}(a|o)\leftarrow$ softmax$\left(\frac{1}{\tau_{q}}\check{Q}_{\theta^{i}_{k+1}}(o,\cdot)\right)(a)$
@@
23:        $\forall i:J^{i}_{t}\leftarrow\{j\in\mathcal{N}:(i,j)\in\mathcal{E}^{comm}_{t}$}
24:        $\forall i:$ Select $\mathrm{adopted}^{i}\sim$ Pr$\left(\mathrm{adopted}^{i}=j\right)$ = $\frac{\exp{(\sigma^{j}_{k+1}}/\tau^{comm}_{k})}{\sum_{x\in J^{i}_{t}}\exp{(\sigma^{x}_{k+1}}/\tau^{comm}_{k})}$ $\forall j\in J^{i}_{t}$
@@
3 Learning and policy improvement
3.1 Q-network and update
Lines 1-14 of our novel Alg. 1 contain the core Q-function/policy update method. Agent $i$ has a neural network parametrised by $\theta^{i}_{k}$ to approximate its Q-function: $\check{Q}_{\theta^{i}_{k}}(o,\cdot)$. The agent’s policy is given by $\pi_{\theta^{i}_{k}}(a|o)=$ softmax$\left(\frac{1}{\tau_{q}}\check{Q}_{\theta^{i}_{k}}(o,\cdot)\right)(a)$. We denote the policy $\pi^{i}_{k}(a|o)$ for simplicity when appropriate. Each agent maintains a buffer (of size $M$) of collected transitions of the form $\left(o^{i}_{t},a^{i}_{t},r^{i}_{t},o^{i}_{t+1}\right)$. At each iteration $k$, they empty their buffer (Line 3) before collecting $M$ new transitions (Lines 4-7); each decentralised agent $i$ then trains its Q-network $\check{Q}_{\theta^{i}_{k}}$ via $L$ training updates as follows (Lines 8-12). For training purposes, $i$ also maintains a target network $\check{Q}_{\theta^{i,^{\prime}}_{k,l}}$ with the same architecture but parameters $\theta^{i,^{\prime}}_{k,l}$ copied from $\theta^{i}_{k,l}$ less regularly than $\theta^{i}_{k,l}$ themselves are updated, i.e. only every $\nu$ learning iterations (Line 11). At each iteration $l$, the agent samples a random batch $B^{i}_{k,l}$ of $|B|$ transitions from its buffer (Line 9), and trains its neural network via stochastic gradient descent to minimise the empirical loss (Def. 8, Line 10). For $cl<0$, $[\cdot]^{0}_{cl}$ is a clipping function used in Munchausen RL to prevent numerical issues if the policy is too close to deterministic, as the log-policy term is otherwise unbounded Vieillard et al. (2020); Wu et al. (2024):
Definition 8 (Empirical loss for Q-network).
This is given by
@@ 3.1 Q-network and update
^$\hat{\mathcal{L}}(\theta,\theta^{\prime})=\frac{1}{|B|}\sum_{transition\in B^{i}_{k,l}}\left|\check{Q}_{\theta^{i}_{k,l}}(o_{t},a_{t})-T\right|^{2},$
where the target $T$ is
$T=r_{t}\;+\;[\tau_{q}\ln\pi_{\theta^{i,^{\prime}}_{k,l}}(a_{t}|o_{t})]^{0}_{cl}\;+$
$\gamma\sum_{a\in\mathcal{A}}\pi_{\theta^{i,^{\prime}}_{k,l}}(a|o_{t+1})\left(\check{Q}_{\theta^{i,^{\prime}}_{k,l}}(o_{t+1},a)-\tau_{q}\ln\pi_{\theta^{i,^{\prime}}_{k,l}}(a|o_{t+1})\right).$
3.2 Communication and adoption of parameters
@@
We use the communication network $\mathcal{G}^{comm}_{t}$ to share two types of information at different points in Alg 1. One is used to improve local estimates of the mean-field distribution (see Sec. 4). The other, described here, is used to privilege the spread of better performing policy updates through the population, allowing faster convergence in this networked case than in the independent and even centralised cases.
@@
We adapt the work in Benjamin and Abate (2023) for the function-approximation case, where in our work agents broadcast the parameters of the Q-network that defines their policy, rather than the Q-function table. At each iteration $k$, after independently updating their Q-network and policy (Lines 3-14), agents approximate the infinite discounted return (Def. 3) of their new policies by collecting rewards for $E$ steps, and assign the finite-step discounted sum to $\sigma^{i}_{k+1}$ (Lines 15-20). They then broadcast their Q-network parameters along with $\sigma^{i}_{k+1}$ (Line 22). Receiving these from neighbours on the network, agents select which set of parameters to adopt by taking a softmax over their own and the received estimate values $\sigma^{j}_{k+1}$ (Lines 23-25). They repeat the process for $C_{p}$ rounds. This allows decentralised agents to adopt policy parameters estimated to perform better than their own, accelerating learning as shown in Sec. 5.
4 Mean-field estimation and communication
@@
We first describe the most general version of our algorithm for decentralised estimation of the empirical categorical mean-field distribution, assuming the more general setting where $\mathcal{G}^{obs}_{t}$ applies (see discussion in Sec. 2.3). We subsequently detail how the algorithm can be made more efficient in environments where the more abstract visibility graph $\mathcal{G}^{vis}_{t}$ applies, as in our experimental settings. In both cases, the algorithm runs to generate the observation object when a step is taken in the main Alg. 1, i.e. to produce $o^{i}_{t}=(s^{i}_{t},\tilde{\hat{\mu}}^{i}_{t})$ for the steps $a^{i}_{t}\sim\pi^{i}_{k}(\cdot|o^{i}_{t})$ in Lines 5, 17 and 26. Both versions of the algorithm are subject to implicit assumptions, which we highlight and discuss methods for addressing in Sup. Mat. E.
4.0.1 Algorithm for the general setting
Algorithm 2 Mean-field estimation and communication in general settings
@@ 4.0.1 Algorithm for the general setting
1:  $\forall i,s:$ Initialise count vector $\hat{\upsilon}^{i}_{t}[s]$ with $\emptyset$
2:  $\forall i:$ $\hat{\upsilon}^{i}_{t}[s^{j}_{t}]$leftarrow $\{ID^{j}\}_{j\in\mathcal{N}:(i,j)\in\mathcal{E}^{obs}_{t}}$
@@
8:  $\forall i:$ $counted\_agents^{i}_{t}\leftarrow\sum_{s\in\mathcal{S}:\hat{\upsilon}^{i}_{t}[s]\neq\emptyset}|\hat{\upsilon}^{i}_{t}[s]|$
9:  $\forall i:$ $uncounted\_agents^{i}_{t}\leftarrow N-counted\_agents^{i}_{t}$
10:  $\forall i,s$ : $\tilde{\hat{\mu}}^{i}_{t}[s]\leftarrow\frac{uncounted\_agents^{i}_{t}}{N\times|\mathcal{S}|}$
11:  $\forall i,s$ where $\hat{\upsilon}^{i}_{t}[s]$ is not $\emptyset$ : $\tilde{\hat{\mu}}^{i}_{t}[s]\leftarrow\tilde{\hat{\mu}}^{i}_{t}[s]+\frac{|\hat{\upsilon}^{i}_{t}[s]|}{N}$
@@
In this setting, our method (Alg. 2) assumes each agent is associated with a unique ID to avoid the same agents being counted multiple times. Each agent maintains a ‘count’ vector ${\hat{\upsilon}}^{i}_{t}$ of length $|\mathcal{S}|$ i.e. of the same shape as the vector denoting the true empirical categorical distribution of agents. Each state position in the vector can hold a list of IDs. Before any actions are taken at each time step $t$, each agent’s count vector ${\hat{\upsilon}}^{i}_{t}$ is initialised as full of $\emptyset$ (‘no count’) markers for each state. Then, for each agent $j$ with which agent $i$ is connected via the observation graph, $i$ places $j$’s unique ID in its count vector in the correct state position. Next, for $C_{e}\geq 0$ communication rounds, agents exchange their local counts with neighbours on the communication network, and merge these counts with their own count vector, filtering out the unique IDs of those that have already been counted. If $C_{e}=0$ then the local count will remain purely independent. By exchanging these partially filled vectors, agents are able to improve their local counts by adding the states of agents that they have not been able to observe directly themselves.
@@
After the $C_{e}$ communication rounds, each state position ${\hat{\upsilon}}^{i}_{t}[s]$ either still maintains the $\emptyset$ marker if no agents have been counted in this state, or contains $x_{s}>0$ unique IDs. The local mean-field estimate $\tilde{\hat{\mu}}^{i}_{t}$ is then obtained from ${\hat{\upsilon}}^{i}_{t}$ as follows. All states that have a count $x_{s}$ have this count converted into the categorical probability $x_{s}/N$ (we assume that agents know the total number of agents in the finite population, even if they cannot observe them all at each $t$). The total number of agents counted in ${\hat{\upsilon}}^{i}_{t}$ is given by $counted\_agents$ = $\sum_{s\in\mathcal{S}}x_{s}$, and the agents that have not been observed are $uncounted\_agents$ = $N$ $counted\_agents$. In this general setting, the unobserved agents are assumed to be uniformly distributed across all the states, so $uncounted\_agents/(N\times|\mathcal{S}|)$ is added to all the values in $\tilde{\hat{\mu}}^{i}_{t}$, replacing the $\emptyset$ marker for states for which no agents have been observed.
4.0.2 Algorithm for visibility-based environments
We give now the differences in our estimation algorithm (Alg. 3, Sup. Mat. A) for the subclass of environments where $\mathcal{G}^{vis}_{t}$ applies in place of $\mathcal{G}^{obs}_{t}$, i.e. the mutual observability of agents depends in turn on the mutual visibility of states. The benefit of $\mathcal{G}^{vis}_{t}$ over $\mathcal{G}^{obs}_{t}$ is that the former allows an agent in state $s$ to obtain a correct, complete count $x_{s^{\prime}}\geq 0$ of all the agents in state $s^{\prime}$, for any state $s^{\prime}$ that is visible to $s$ (note the count may be zero). Unique IDs are thus not required as there is no risk of counting the same agent twice when receiving communicated counts: either all agents in $s^{\prime}$ have been counted, or no count has yet been obtained for $s^{\prime}$. This simplifies the algorithm and helps preserve agent privacy.
Secondly, uncounted agents cannot be in states for which a count has already been obtained, since the count is complete and correct, even if the count is $x_{s^{\prime}}=0$. Therefore after the $C_{e}$ communication rounds, the $uncounted\_agents$ proportion needs to be uniformly distributed only across the positions in the vector that still have the $\emptyset$ marker, and not across all states as in the general setting. This makes the estimation more accurate in this special setting.
5 Theoretical results
To demonstrate the benefits of the networked architecture by comparison, we also consider (theoretically here and experimentally in Sec. 6) the results of modified versions of our algorithm for centralised and independent learners. In the centralised setting, the Q-network updates of arbitrary agent $i=1$ are automatically pushed to all other agents, and the true global mean-field distribution is always used in place of the local estimate i.e. $\tilde{\hat{\mu}}^{i}_{t}$ = ${\hat{\mu}}_{t}$. In the independent case, there are no links in $\mathcal{G}^{comm}_{t}$ or $\mathcal{G}^{vis}_{t}$, i.e. $\mathcal{E}^{comm}_{t}=\mathcal{E}^{vis}_{t}=\emptyset$.
Networked agents often learn faster than centralised ones in our experiments; we justify theoretically this possibly counterintuitive result here. We first make two strong assumptions that give conditions under which networked agents do outperform centralised ones. The fact that these strong assumptions do not always hold in reality explains why networked agents may not always outperform centralised ones.
Recall that at each iteration $k$ of Alg. 1, after independently updating their policies in Line 14, the population has the policies $\pi^{i}_{k+1}$}${}_{i=1}^{N}$. There is randomness in these independent policy updates, stemming from the random sampling of each agent’s independently collected buffer. In Lines 15-20, agents approximate the infinite discounted returns {$V(\pi^{i}_{k+1},I(\pi^{i}_{k+1}))\}_{i=1}^{N}$ (Def. 3) of their updated policies by computing $\sigma^{i}_{k+1}$}${}_{i=1}^{N}$: the $E$-step discounted return with respect to the empirical mean field generated when agents follow policies $\pi^{i}_{k+1}$}${}_{i=1}^{N}$ (i.e. they do not at this stage all follow a single identical policy). We can characterise the approximation as $\sigma^{i}_{k+1}$}${}_{i=1}^{N}$ = $\widehat{V}(\pi^{i}_{k+1},I(\pi^{i}_{k+1}))\}_{i=1}^{N}$.
Assumption 1.
Assume that $\sigma^{i}_{k+1}$}${}_{i=1}^{N}$ are sufficiently good approximations so as to respect the ordering of the true values {$V(\pi^{i}_{k+1},I(\pi^{i}_{k+1}))\}_{i=1}^{N}$, i.e. $\forall i,j\in\{1,\dots,N\}$: $\sigma^{i}_{k+1}>\sigma^{j}_{k+1}\iff V(\pi^{i}_{k+1},I(\pi^{i}_{k+1}))>V(\pi^{j}_{k+1},I(\pi^{j}_{k+1})).$
Assumption 2.
Assume that after the $C_{p}$ rounds in Lines 21-27 in which agents exchange and adopt policies from neighbours, the population is left with a single policy such that $\forall i,j\in\{1,\dots,N\}$ $\pi^{i}_{k+1}=\pi^{j}_{k+1}$. [^2] Most simply we can think of Ass. 2 holding if 1) $\tau^{comm}_{k}\rightarrow 0\;\forall k$ such that the softmax essentially becomes a max function, and 2) the communication network $\mathcal{G}^{comm}_{t}$ is static and connected during the $C_{p}$ communication rounds, where $C_{p}$ is larger than the network diameter $d_{\mathcal{G}^{comm}_{t}}$ Under these conditions, previous results on max-consensus algorithms show that all agents in the network will converge on the highest $\sigma^{max}_{k+1}$ value (and hence the unique associated $\pi^{max}_{k+1}$ within a number of rounds equal to the diameter $d_{\mathcal{G}^{comm}_{t}}$ Nejad et al. (2009); Benjamin and Abate (2023). However, policy consensus as in Ass. 2 might be achieved even outside of these conditions, including if the network is dynamic and not connected at every step, given appropriate values for $C_{p}$ and $\tau^{comm}_{k+1}\in\mathbb{R}_{>0}$.
Call the network consensus policy $\pi^{\mathrm{net}}_{k+1}$, and its associated finitely approximated return $\sigma^{\mathrm{net}}_{k+1}$. Recall that the centralised case is where the Q-network update of arbitrary agent $i=1$ is automatically pushed to all the others instead of the policy evaluation and exchange in Lines 15-27; this is equivalent to a networked case where policy consensus is reached on a random one of the policies $\pi^{i}_{k+1}$}${}_{i=1}^{N}$. Call this policy arbitrarily given to the whole population $\pi^{\mathrm{cent}}_{k+1}$, and its associated finitely approximated return $\sigma^{\mathrm{cent}}_{k+1}$.
Theorem 1.
Given Ass. 1 and 2, $\mathbb{E}[V(\pi^{\mathrm{net}}_{k+1},I(\pi^{\mathrm{net}}_{k+1}))]>\mathbb{E}[V(\pi^{\mathrm{cent}}_{k+1},I(\pi^{\mathrm{cent}}_{k+1}))]$. Thus in expectation networked agents will increase their returns faster than centralised ones.
Proof.
Recall that before the communication rounds in Line 21 (Alg. 1), the randomly updated policies $\pi^{i}_{k+1}$}${}_{i=1}^{N}$ have associated approximated returns $\sigma^{i}_{k+1}$}${}_{i=1}^{N}$. Denote the mean and maximum of this set $\sigma^{\mathrm{mean}}_{k+1}$ and $\sigma^{\mathrm{max}}_{k+1}$ respectively. Since $\pi^{\mathrm{cent}}_{k+1}$ is chosen arbitrarily from $\pi^{i}_{k+1}$}${}_{i=1}^{N}$, it will obey $\mathbb{E}[\sigma^{\mathrm{cent}}_{k+1}]=\sigma^{\mathrm{mean}}_{k+1}$ $\forall k$, though there will be high variance. Conversely, the softmax adoption probability (Line 24, Alg. 1) for the networked case means by definition that policies with higher $\sigma^{i}_{k+1}$ are more likely to be adopted at each communication round. Thus the $\pi^{\mathrm{net}}_{k+1}$ that gets adopted by the whole networked population will obey $\mathbb{E}[\sigma^{\mathrm{net}}_{k+1}]>\sigma^{\mathrm{mean}}_{k+1}$ (if $\tau^{comm}_{k+1}\rightarrow 0$, it will obey $\mathbb{E}[\sigma^{\mathrm{net}}_{k+1}]=\sigma^{\mathrm{max}}_{k+1}$ $\forall k$). As such, $\mathbb{E}[\sigma^{\mathrm{net}}_{k+1}]>\mathbb{E}[\sigma^{\mathrm{cent}}_{k+1}]$, which by Ass. 1 implies the result. ∎
The adoption scheme in Line 24 biases the spread of policies towards those estimated to be better, which, given sufficiently good approximations (Ass. 1), results in higher discounted returns in practice. By choosing updates in a more principled way, networked agents learn faster than the centralised case that adopts updates regardless of quality. This intuition applies even if we loosen Ass. 2 that the networked population converges on a single consensus policy within the $C_{p}$ communication rounds. Similar logic can also be applied to understand why networked agents outperform entirely independent ones, combined with the fact that divergence between policies in the independent case worsens sample complexity over the networked and centralised cases by biasing approximations of the Q-function Yardim et al. (2023); Benjamin and Abate (2023).
Significantly, the communication scheme not only allows us to avoid the undesirable assumption of a central learner, but even to outperform it. Moreover, the benefit of networked communication over centralised learning is even greater in our function approximation setting than in the tabular case of Benjamin and Abate (2023), perhaps due to greater variance in the quality of Q-function estimates in our case.
6 Experiments
We provide two sets of experiments. The first set showcases that our function-approximation algorithm (Alg. 1) can scale to large state spaces for population-independent policies, and that in such settings networked, communicating agents can outperform purely-independent and even centralised agents, and do so by an even greater margin than in the tabular settings from Benjamin and Abate (2023). The second set (given in Sup. Mat. C.4) demonstrates that Alg. 1 can handle population-dependent policies, as well as the ability of Alg. 3 to practically estimate the mean-field distribution locally.
For the types of game used in our demonstrations we follow the gold standard in prior MFG works, i.e. grid-world environments where agents can move in the four cardinal directions or remain in place Laurière (2021); Laurière et al. (2022b); Zaman et al. (2023); Algumaei et al. (2023); Cui et al. (2023); Benjamin and Abate (2023); Wu et al. (2024). We present results from five tasks defined by the agents’ reward/transition functions, all but one of which are coordination tasks. The first two tasks are those used with population-independent policies in Benjamin and Abate (2023), but while they show results for an 8x8 and a ‘larger’ 16x16 grid, our results are for 100x100 and 50x50 grids:
See Sup. Mat. C.1 for a full technical description of these tasks, as well as descriptions of additional more complex tasks. In these spatial environments, both the communication network $\mathcal{G}^{comm}_{t}$ and the visibility graph $\mathcal{G}^{vis}_{t}$ are determined by the physical distance from agent $i$; we show plots for various radii, expressed as fractions of the maximum possible distance (the grid’s diagonal length). We discuss hyperparameters in Sup. Mat. C.3.
Figure 1: ‘Target agreement’, pop.-independent, 100x100 grid.
Figure 2: ‘Cluster’, pop.-independent, 100x100 grid.
We evaluate our experiments via two metrics. Exploitability is the most common metric in works on MFGs, and is a measure of proximity to the MFNE. It quantifies how much a best-responding agent can benefit by deviating from the set of policies that generate the current mean-field distribution, with a decreasing exploitability meaning the population is closer to the MFNE. However, there are several issues with this metric in our setting, particularly for our coordination tasks where competitive agents benefit from aligning behaviours, such that it may give limited or noisy information (discussed further in Sup. Mat. C.2.1). We thus also give a second metric, as in Benjamin and Abate (2023): the population’s average discounted return. This allows us to compare how quickly agents are learning to increase their returns, even when ‘exploitability’ gives us limited ability to distinguish between the desirability of the MFNEs to which populations converge.
6.1 Results and discussion
Figs. 1 and 2 illustrate that introducing function approximation to algorithms in this setting allows them to converge within a practical number of iterations ($k\ll 100$), even for large state spaces (100x100 grids). By contrast, the tabular algorithms in Benjamin and Abate (2023) appear only just to converge by $k=200$ for the same tasks for the larger of their two grids, which is only 16x16.
In Figs. 1 and 2, networked agents all significantly outperform both centralised and independent agents in term of average return, despite the centralised agents appearing to have similar exploitability, and the independent agents having similar or slightly lower exploitability. This is because independent agents (and also the centralised ones in Fig. 1) hardly improve their policies at all, such that there is little a deviating agent can do to increase its return in these coordination games, meaning exploitability appears low, despite this being an undesirable equilibrium (see Sup. Mat. C.2 for further discussion on the limited information provided by the exploitation metric). The fact that the networked agents nevertheless significantly outperform the other architectures in terms of average return indicates that communication helps agents to find substantially ‘preferable’ equilibria. Moreover, the margin by which the networked agents can outperform the centralised agents is much greater than in Benjamin and Abate (2023), showing that the benefits of the communication scheme are even greater in non-tabular settings. See Sup. Mat. C.4 for further experiments with large state spaces, as well as others demonstrating the ability of our algorithm to handle two more complex tasks, using population-dependent policies and estimated mean-field observations.
7 Conclusion
We novelly contributed function approximation to the online-learning setting for empirical MFGs, and also contributed two novel algorithms for locally estimating the empirical mean field for population-dependent policies. We have justified theoretically why our networked communication algorithm is able to learn faster than both centralised and independent agents in this function approximation setting, and demonstrated empirically the ability of our algorithms to handle large state spaces and estimate the mean field. Limitations and ways to extend our algorithms are in Sup. Mat. E.
Supplementary material to ‘Networked Communication for Mean-Field Games with Function Approximation and Empirical Mean-Field Estimation’
Appendix A Mean-field estimation algorithm for visibility-based environments
Algorithm 3 Mean-field estimation and communication for environments with $\mathcal{G}^{vis}_{t}$
@@ Appendix A Mean-field estimation algorithm for visibility-based environments
1:  $\forall i,s:$ Initialise count vector $\hat{\upsilon}^{i}_{t}[s]$ with $\emptyset$
2:  $\forall i$ and $\forall s^{\prime}\in\mathcal{S^{\prime}}:(s^{i}_{t},s^{\prime})\in\mathcal{E}^{vis}_{t}$ : $\hat{\upsilon}^{i}_{t}[s^{\prime}]$leftarrow $\sum_{j\in 1,\dots,N:s^{j}_{t}=s^{\prime}}1$

### v3

@@ 3.3. Networks
Definition 3.6 (Time-varying network).
The network ($\mathcal{G}_{t}$)t≥0 is given by $\mathcal{G}_{t}$ = ($\mathcal{N},\mathcal{E}_{t}$), where $\mathcal{N}$ is the set of vertices representing agents $i\in\{1,\dots,N\}$, and the edge set $\mathcal{E}_{t}$ $\subseteq$ {(i,j) : i,j $\in$ $\mathcal{N}$, i $\neq$ j} is the set of undirected links present at time t. A network’s diameter $d_{\mathcal{G}_{t}}$ is the maximum of the shortest path length between any pair of nodes.
We also define an alternative version of the observation graph that is useful in a specific subclass of environments, which can most intuitively be thought of as those where agents’ states are positions in physical space. When this is the case, we usually think of agents’ ability to observe each other as depending more abstractly on whether states are visible to each other. This visibility graph is:
Definition 3.7 (Time-varying state-visibility graph).
The state visibility graph ($\mathcal{G}^{vis}_{t}$)t≥0 is given by $\mathcal{G}^{vis}_{t}$ = ($\mathcal{S^{\prime}},\mathcal{E}^{vis}_{t}$), where $\mathcal{S^{\prime}}$ is the set of vertices representing the environment states $\mathcal{S}$, and the edge set $\mathcal{E}^{vis}_{t}$ $\subseteq$ {(m,n) : m,n $\in$ $\mathcal{S^{\prime}}$} is the set of undirected links present at time t, indicating which states are visible to each other.
We say an agent in $s$ can obtain a count of the number of agents in $s^{\prime}$ if $s^{\prime}$ is visible to $s$. The benefit of this graph $\mathcal{G}^{vis}_{t}$ over $\mathcal{G}^{obs}_{t}$ is that there is mutual exclusivity: either an agent in state $s$ is able to obtain a total count of all of the agents in state $s^{\prime}$ (if $s^{\prime}$ is visible to $s$), or it cannot obtain information about any agent in state $s^{\prime}$ (if those states are not visible to each other). Additionally, this graph permits an agent in state $s$ to observe that there are no agents in state $s^{\prime}$ as long as $s^{\prime}$ is visible to $s$. These benefits are not available if the observability graph is defined strictly between agents as in $\mathcal{G}^{obs}_{t}$, such that using $\mathcal{G}^{vis}_{t}$ facilitates more efficient estimation of the global mean-field distribution from local information in settings where $\mathcal{G}^{vis}_{t}$ is applicable (see Sec. 5).
Algorithm 1 Networked learning with function approximation
0: loop parameters $K,M,L,E,C_{p}$; learning parameters $\gamma,\tau_{q},|B|,cl,\nu$, $\{\tau^{comm}_{k}\}_{k\in\{0,\dots,K-1\}}$; initial states $s^{i}_{0}$$_{i=1}^{N}$; $t\leftarrow 0$
@@
10:   Update $\theta$ to minimise $\hat{\mathcal{L}}(\theta,\theta^{\prime})$ as in Def. 4.1
@@
[…]
@@
23:   $\forall i:J^{i}_{t}\leftarrow i\cup\{j\in\mathcal{N}:(i,j)\in\mathcal{E}^{comm}_{t}$}
[…]
@@
4. Learning and policy improvement
4.1. Q-network and update
[…]
Definition 4.1 (Empirical loss for Q-network).
For the target $T$:
$T=r_{t}$ $+[\tau_{q}\ln\pi_{\theta^{i,^{\prime}}_{k,l}}(a_{t}|o_{t})]^{0}_{cl}$
$+\gamma\sum_{a\in\mathcal{A}}\pi_{\theta^{i,^{\prime}}_{k,l}}(a|o_{t+1})\left(\check{Q}_{\theta^{i,^{\prime}}_{k,l}}(o_{t+1},a)-\tau_{q}\ln\pi_{\theta^{i,^{\prime}}_{k,l}}(a|o_{t+1})\right),$
@@ 4.1. Q-network and update
$\text{the loss is:}\;\;\;\;\;\;\;\;\;\;\hat{\mathcal{L}}(\theta,\theta^{\prime})=\frac{1}{|B|}\sum_{transition\in B^{i}_{k,l}}\left|\check{Q}_{\theta^{i}_{k,l}}(o_{t},a_{t})-T\right|^{2}.$
4.2. Communication and adoption of parameters
@@
We use the communication network $\mathcal{G}^{comm}_{t}$ to share two types of information at different points in Alg 1. One is used to improve local estimates of the mean field (Sec. 5). The other, described here, is used to privilege the spread of better performing policy updates through the population, allowing faster convergence in this networked case than in the independent and even central-agent cases.
@@
We adapt benjamin2024networked for the function-approximation case, where in our work agents broadcast the parameters of the Q-network that defines their policy, rather than the Q-function table. At each iteration $k$, after independently updating their Q-network and policy (Lines 3-14), agents estimate the infinite discounted return (Def. 3.3) of their new policies by collecting rewards for $E$ steps, and assign the finite-step discounted sum to $\sigma^{i}_{k+1}$ (Lines 15-20). They then broadcast their Q-network parameters along with $\sigma^{i}_{k+1}$ (Line 22). Receiving these from neighbours on the network, agents select which set of parameters to adopt by taking a softmax over their own and the received estimate values $\sigma^{j}_{k+1}$ (Lines 23-25). They repeat the process for $C_{p}$ rounds. This allows decentralised agents to adopt policy parameters estimated to perform better than their own, accelerating learning as shown in Sec. 6.
5. Mean-field estimation and communication
@@
We now give our algorithms for decentralised estimation of the empirical categorical mean-field distribution. We first describe the general version, assuming the more general setting where $\mathcal{G}^{obs}_{t}$ applies (see discussion in Sec. 3.3). We subsequently detail how the algorithm can be made more efficient in environments where the more abstract visibility graph $\mathcal{G}^{vis}_{t}$ applies, as in our experimental settings. In both cases, the algorithm runs to generate the observation object when a step is taken in the main Alg. 1, i.e. to produce $o^{i}_{t}=(s^{i}_{t},\tilde{\hat{\mu}}^{i}_{t})$ for the steps $a^{i}_{t}\sim\pi^{i}_{k}(\cdot|o^{i}_{t})$ in Lines 5, 17 and 26. Note that if $\mathcal{G}^{obs}_{t}$$\mathcal{G}^{vis}_{t}$ are fully connected, all agents’ estimated mean-field observations will be equivalent to the true categorical distribution. Both versions of the algorithm are subject to implicit assumptions, which we discuss methods for addressing as future work in Appx. C.
5.1. Algorithm for the general setting
Algorithm 2 Mean-field estimation in general settings
@@ 5.1. Algorithm for the general setting
[…]
@@
[…]
@@
In this setting, our method (Alg. 2) assumes each agent is associated with a unique ID to avoid the same agents being counted multiple times. Each agent maintains a ‘count’ vector ${\hat{\upsilon}}^{i}_{t}$ of length $|\mathcal{S}|$ i.e. of the same shape as the vector denoting the true empirical categorical distribution of agents. Each state position in the vector can hold a list of IDs. Before any actions are taken at each time step $t$, each agent’s count vector ${\hat{\upsilon}}^{i}_{t}$ is initialised as full of $\emptyset$ (‘no count’) markers for each state (Line 1). Then, for each agent $j$ with which agent $i$ is connected via the observation graph, $i$ places $j$’s unique ID in its count vector in the correct state position (Line 2). Next, for $C_{e}\geq 0$ communication rounds, agents exchange their local counts with neighbours on the communication network (Line 4), and merge these counts with their own count vector, filtering out the unique IDs of those that have already been counted (Line 6). If $C_{e}=0$ then the local count will remain purely independent. By exchanging these partially filled vectors, agents are able to improve their local counts by adding the states of agents that they have not been able to observe directly themselves.
@@
After the $C_{e}$ communication rounds, each state position ${\hat{\upsilon}}^{i}_{t}[s]$ either still maintains the $\emptyset$ marker if no agents have been counted in this state, or contains $x_{s}>0$ unique IDs. The local mean-field estimate $\tilde{\hat{\mu}}^{i}_{t}$ is then obtained from ${\hat{\upsilon}}^{i}_{t}$ as follows. All states that have a count $x_{s}$ have this count converted into the categorical probability $x_{s}/N$ (we assume that agents know the total number of agents in the finite population, even if they cannot observe them all at each $t$) (Line 11). The total number of agents counted in ${\hat{\upsilon}}^{i}_{t}$ is given by counted_agentscounted\_agents = $\sum_{s\in\mathcal{S}}x_{s}$, and the agents that have not been observed are uncounted_agentsuncounted\_agents = $N$ - counted_agentscounted\_agents. In this general setting, the unobserved agents are assumed to be uniformly distributed across all the states, so $uncounted\_agents/(N\times|\mathcal{S}|)$ is added to all the values in $\tilde{\hat{\mu}}^{i}_{t}$, replacing the $\emptyset$ marker for states for which no agents have been observed (Line 10).
Algorithm 3 Mean-field estimation for environments with $\mathcal{G}^{vis}_{t}$
@@
[…]